# Wavelet-based Multi-View Fusion of 4D Radar Tensor and Camera for Robust 3D Object Detection

### v1

@@ Abstract
4D millimeter-wave (mmWave) radar has been widely adopted in autonomous driving and robot perception due to its low cost and all-weather robustness. However, its inherent sparsity and limited semantic richness significantly constrain perception capability. Recently, fusing camera data with 4D radar has emerged as a promising cost effective solution, by exploiting the complementary strengths of the two modalities. Nevertheless, point-cloud-based radar often suffer from information loss introduced by multi-stage signal processing, while directly utilizing raw 4D radar data incurs prohibitive computational costs. To address these challenges, we propose WRCFormer, a novel 3D object detection framework that fuses raw radar cubes with camera inputs via multi-view representations of the decoupled radar cube. Specifically, we design a Wavelet Attention Module as the basic module of wavelet-based Feature Pyramid Network (FPN) to enhance the representation of sparse radar signals and image data. We further introduce a two-stage query-based, modality-agnostic fusion mechanism termed Geometry-guided Progressive Fusion to efficiently integrate multi-view features from both modalities. Extensive experiments demonstrate that WRCFormer achieves state-of-the-art performance on the K-Radar benchmarks, surpassing the best model by approximately 2.4% in all scenarios and 1.6% in the sleet scenario, highlighting its robustness under adverse weather conditions.
@@ I Introduction
Robust perception is essential for autonomous driving and robotic navigation. Modern perception systems commonly rely on sensor suites, including cameras, LiDAR, and radar [1]. Cameras capture rich information such as color and texture, while LiDAR provides detailed three-dimensional (3D) geometric representations of the environment [2, 3, 4]. However, neither cameras nor LiDAR can directly measure object velocity, and both sensors are highly sensitive to illumination variations, particularly under adverse weather conditions such as rain, fog, or snow. In contrast, recent advances in 4D radar has shown the enhanced ability to capture vertical information compared to conventional 3D radars [5]. Despite this progress, many existing 4D radar perception methods based on point cloud representations remain constrained by sparsity [6, 7, 8, 9, 10]. Fortunately, radar and camera exhibit strong complementarity in object representation [11, 12, 13, 14, 15, 16], making a promising low-cost and efficient solution for robust perception.
@@
Conventional convolutions often struggle to capture the multi-scale and oscillatory patterns in radar tensors, leading to blurred boundaries and insufficient noise suppression. In contrast, wavelet-based modules offer joint spatial-frequency representations that strengthen foreground signatures while effectively suppressing background clutter. To leverage the intrinsic spectral-spatial coupling in radar data, we design a Wavelet Attention Mixture-of-Experts (WA-MoE) module as the core of a sensor-oriented Feature Pyramid Network (FPN). Despite the complementarity of radar and camera, current fusion methods often incur high computational costs and face challenges in achieving fine-grained cross-modal alignment. To overcome these limitations, we introduce a Geometry-guided Progressive Fusion (GPF) strategy that uses geometric priors to progressively align radar and visual features, achieving higher detection accuracy with notably reduced computational overhead.
@@ II Related Works
Millimeter-wave (mmWave) radar serves as an all-weather perception sensor capable of measuring the 3D position, velocity, and RCS of objects. Recent advances in 4D mmWave radar, which provides denser point clouds than conventional 3D radar, have enabled standalone 3D object detection through deep learning models based on 3D voxel representations [6, 7, 25, 26, 27, 9]. Meanwhile, radar-camera fusion has shown promising performance by exploiting the complementary strengths of both modalities in a cost-effective manner [5, 21]. Early fusion approaches typically combined radar point clouds with camera images for downstream perception tasks such as object detection [12, 15, 11, 9, 28]. Another line of work explores radar-LiDAR fusion, where high-resolution geometric features from LiDAR complements radar features for improved detection performance [29, 30, 31, 32]. Despite their effectiveness, point cloud representations remain inherently sparse and suffer from significant information loss due to multi-stage radar signal processing. To overcome this limitation, we propose an efficient fusion strategy that decomposes the raw 4D radar tensor into multi-view representations for integration with visual features, which preserves richer radar information while enabling robust radar-camera fusion for 3D object detection.
II-B 3D Perception with Pre-CFAR Radar Data
@@
Previous studies, including approaches based on radar point clouds and tensor representations, have shown that radar can enhance 3D object detection [33] and depth estimation [34]. However, conventional 3D radar lacks elevation resolution, limiting its ability to capture rich semantic and geometric cues, which constrains perception accuracy [21]. This limitation has motivated a shift toward utilizing 4D radar tensors, which preserve the complete signal structure without projection-induced information loss. Research in this direction often employs pre-CFAR or raw radar data representations. Early efforts focused on fusing radar and camera data at the feature level for vehicle detection [35], while subsequent works directly processed 3D radar tensors (range-azimuth-Doppler) using deep learning for improved detection [36]. To better exploit radar characteristics, researchers have proposed specialized network architectures, including RAMP-CNN for enhanced object recognition [37], RODNet for cross-supervised radar detection fused with camera localization [38], and RaDDet for detecting dynamic road users [39]. Further advancements include Radatron, which uses cascaded multi-resolution radar inputs [40], and transformer-based models like T-FFTRadNet operating on raw ADC signals [41]. Researchers have also addressed efficient online detection using recurrent CNNs [42] and sparse networks like SparseRadNet [43]. More recently, methods including RTNH+ [44] and SpikingRTNH [45] have further advanced 4D radar object detection through improved pre-processing strategies and spiking neural network architectures. Several datasets support this line of research. RADIal [46] provides ADC signals, range-azimuth maps, and point clouds with segmentation annotations. Other datasets, including VoD [47], TJ4DRadSet [48], Dual Radar [8], MAN TruckScenes [49], and OmniHD-Scenes [50], primarily offer radar point clouds. In contrast, our work builds upon the K-Radar dataset [20], which is currently the only public 4D radar tensor dataset featuring multiple object types across diverse environments alongside corresponding 3D bounding box annotations. This unique combination enables a comprehensive evaluation of tensor-based 3D detection methods.
@@ III Method
We propose WRCFormer, a novel architecture designed for robust sensor fusion that processes inputs from a monocular camera and a 4D radar tensor as Fig. 3. The framework begins with a parallel feature extraction pipeline, in which an Image Encoder and a Feature Pyramid Network (FPN) generate multi-scale visual features. Concurrently, to reduce the resolution mismatch between image and radar data, we adopt 4D radar tensor (cube) and decompose it into Elevation-Azimuth (EA) and Range-Azimuth (RA) representations. These radar maps are fed into specialized encoders and an FPN based on our proposed Wavelet Attention Mixture-of-Experts (WA-MoE) module, which decomposes into complementary high- and low-frequency components and adaptively fuse them. The Geometry-guided Progressive Fusion (GPF) module serves as the core fusion component of WRCFormer and performs a two-stage fusion process. In the first stage, Geometry-driven Semantic Alignment employs a Fast Cross Attention mechanism aligns the semantic image map with the geometric EA radar map. In the second stage, Range-aware Geometric Refinement, further refines the radar-aligned visual features, using Deformable Attention Queries that are precisely guided by the radar RA map. Finally, the framework forwards the resulting fused features to a detection head, which applies an iterative refinement strategy to progressively improve predictions and produce the final detection results.
@@ III-C Wavelet Attention Mixture-of-Experts
Conventional convolution-based FPNs rely on local receptive fields for downsampling and feature fusion, which essentially perform localized weighted averaging. However, the 4D radar tensor constitutes a high-noise, multi-scale signalthat exhibits strong spectral–spatial coupling structure. To better capture this structure, we design a Haar wavelet–based neural network for frequency-domain decomposition, motivated by the following considerations. (1) Noise components predominantly concentrate in the high-frequency sub-bands, whereas reflection peaks remain preserved in the low-frequency subbands. This property naturally facilitates target-clutter separation; (2) Radar sensing fundamentally operates in the frequency domain via Doppler and delay sampling, and wavelet tranform aligns well with the underlying radar scattering mechanism; (3) The introduction of a gated MoE is functionally equivalent to an adaptive filter bank, where the optimal spectral weights are estimated online for each channel and spatial location. This design enables data-driven spectrum shaping and offers greater flexible than manually designed band-pass filters, while providing finer granularity than statically weighted Wavelet-CNNs; (4) For visual inputs, the high-frequency components of wavelets explicitly capture object edges, which further facilitate cross-modal alignment during the subsequent fusion process.
@@
where $\oplus$ denotes concatenation operation and $F_{1}$ is the first-order wavelet domain encoding result. This branch directly learns and extracts spatial correlations within each frequency sub-band from the first-level wavelet decomposition. As a result, it captures explicit and prominent spatio-frequency patterns, serving as a “shallow” feature extractor in the wavelet domain.
@@
where $\mathtt{g_{i}}(F_{\text{fused}})$ represents the sparse gating function, which dynamically computes a weighted combination of experts $\mathtt{E_{i}(\cdot)}$ for a given input. Each expert is implemented as a lightweight Feed-Forward Network (FFN) using the depthwise-separable convolutions and specializes in different patterns.
@@ III-D Geometry-guided Progressive Fusion
While radar-camera fusion has demonstrated strong potential under adverse conditions, existing fusion paradigms still face critical challenges. On the one hand, the direct processing of 4D radar tensors incurs prohibitive computational overhead. On the other hand, existing fusion strategies often overlook the progressive nature of geometric priors, leading to suboptimal semantic–geometric alignment between radar signals and image features. These limitations motivate us to design a more efficient and structured fusion mechanism. To this end, we propose Geometry-guided Progressive Fusion (GPF), a two-stage fusion strategy that leverages the complementary strengths of radar geometry and image semantics in a coarse-to-fine manner, which includes Geometry-driven Semantic Alignment (GSA, Stage 1) and Range-aware Geometric Refinement (RGR, Stage 2).
@@
After that, we apply adaptive max-pooling to the EA query $Q\in\mathbb{R}^{N\times d}$ to obtain a compact representation $A\in\mathbb{R}^{n\times d}$, where $n\ll N$. Compared with adaptive average pooling, adaptive max pooling preserves the strongest signal responses in the EA features while effectively suppressing background interference. This operation preserves the most discriminative geometric responses while significantly reducing the sequence length.
@@
Table II: Overall comparison of 3D object detection on K-Radar (v1.0 version) with a threshold of 0.3.
@@ IV-B Quantitative Results
[…]
@@ V Conclusion
In this paper, we introduce WRCFormer, a robust 3D object detector that effectively fuses a monocular camera with raw 4D radar tensors. The proposed method introduces two key innovations: a Wavelet Attention Mixture-of-Experts module that captures joint spatial-frequency representations to enhance sparse radar signatures, and a two-stage Geometry-guided Progressive Fusion mechanism that enables efficient and semantically-aligned cross-modal fusion integration. Extensive experiments on the K-Radar benchmark demonstrate that WRCFormer sets a new state-of-the-art performance, with particularly strong robustness under adverse weather conditions where camera-based perception often degrades. These results underscores the significant advantage of directly leveraging rich 4D radar tensor, paired with specialized neural architectures, for building reliable, all-weather perception systems in autonomous driving.

### v2

@@ Abstract
4D millimeter-wave (mmWave) radar has been widely adopted in autonomous driving and robot perception due to its low cost and all-weather robustness. However, point-cloud-based radar representations suffer from information loss due to multi-stage signal processing, while directly utilizing raw 4D radar tensors incurs prohibitive computational costs. To address these challenges, we propose WRCFormer, a novel 3D object detection framework that efficiently fuses raw 4D radar cubes with camera images via decoupled multi-view radar representations. Our approach introduces two key components: (1) A Wavelet Attention Module embedded in a wavelet-based Feature Pyramid Network (FPN), which enhances the representation of sparse radar signals and image data by capturing joint spatial-frequency features, thereby mitigating information loss while maintaining computational efficiency. (2) A Geometry-guided Progressive Fusion mechanism, a two-stage query-based fusion strategy that progressively aligns multi-view radar and visual features through geometric priors, enabling modality-agnostic and efficient integration without overwhelming computational overhead. Extensive experiments on the K-Radar benchmark show that WRCFormer achieves state-of-the-art performance, surpassing the best existing model by approximately 2.4% in all scenarios and 1.6% in sleet conditions, demonstrating strong robustness in adverse weather.
@@ I Introduction
Robust perception is essential for autonomous driving and robotic navigation. Modern perception systems commonly rely on sensor suites, including cameras, LiDAR, and radar [27]. Cameras capture rich information such as color and texture, while LiDAR provides detailed three-dimensional (3D) geometric representations of the environment [50, 47, 15]. However, neither cameras nor LiDAR can directly measure object velocity, and both sensors are highly sensitive to illumination variations, particularly under adverse weather conditions such as rain, fog, or snow. In contrast, recent advances in 4D radar has shown enhanced abilities to capture vertical information compared to conventional 2D radars [43]. Despite this progress, many existing 4D radar perception methods based on point cloud representations remain constrained by sparsity, i.e., radar point clouds are notably sparse in 3D space compared to dense LiDAR scans [13, 10, 48, 2, 39]. Fortunately, radar and camera exhibit strong complementarity in object representation [51, 37, 38, 9, 1, 52], making them a promising low-cost and efficient solution for robust perception.
@@
Conventional convolution, while effective in many vision tasks, often struggle to capture the multi-scale and oscillatory patterns inherent in raw radar tensors. This limitation stems from their fixed local receptive fields and isotropy, which can blur subtle target boundaries and inadequately suppress spatially distributed clutter noise in radar data [46]. In contrast, wavelet transforms provide a natural and mathematically grounded framework for analyzing signals with joint localization in space and frequency. This property is particularly aligned with the physical nature of radar sensing: scattering events manifest as localized energy concentrations across specific frequency bands (e.g., Doppler bins), while noise and clutter often spread across scales. By deploying learnable wavelet-based modules, our method explicitly decomposes the radar signal into multi-resolution sub-bands. This allows the network to strengthen foreground signatures by concentrating on relevant frequency components while effectively suppressing background clutter by attenuating high-frequency noise and diffuse interference, directly addressing the limitations of conventional convolutions for radar tensor representation.
@@ II Related Works
Millimeter-wave (mmWave) radar serves as an all-weather perception sensor capable of measuring the 3D position, velocity, and RCS of objects. Recent advances in 4D mmWave radar, which provides denser point clouds than conventional 2D radar, have enabled standalone 3D object detection through deep learning models based on 3D voxel representations [13, 10, 19, 42, 35, 2]. Meanwhile, radar-camera fusion has shown promising performance by exploiting the complementary strengths of both modalities in a cost-effective manner [43, 44]. Early fusion approaches typically combined radar point clouds with camera images for downstream perception tasks such as object detection [37, 1, 51, 2, 49]. Another line of work explores radar-LiDAR fusion, where high-resolution geometric features from LiDAR complements radar features for improved detection performance [33, 41, 31, 18]. Despite their effectiveness, point cloud representations remain inherently sparse and suffer from significant information loss due to multi-stage radar signal processing. To overcome this limitation, we propose an efficient fusion strategy that decomposes the raw 4D radar tensor into multi-view representations for integration with visual features, which preserves richer radar information while enabling robust radar-camera fusion for 3D object detection.
II-B 3D Object Detection with Pre-CFAR Radar Data
@@
Previous studies, including approaches based on radar point clouds and tensor representations, have shown that radar can enhance 2D object detection [20] and depth estimation [32]. However, conventional 2D radar lacks elevation resolution, limiting its ability to capture rich semantic and geometric cues, which constrains perception accuracy [44]. This limitation has motivated a shift toward utilizing 4D radar tensors, which preserves the complete signal structure without projection-induced information loss. Research in this direction often employs pre-CFAR or raw radar data representations. Early efforts focus on fusing radar and camera data at the feature level for vehicle detection [12], while subsequent works directly processed 3D radar tensors (range-azimuth-Doppler) using deep learning for improved detection [17]. To better exploit radar characteristics, researchers have proposed specialized network architectures, including RAMP-CNN for enhanced object recognition [7], RODNet for cross-supervised radar detection fused with camera localization [34], and RaDDet for detecting dynamic road users [45]. Further advancements include Radatron, which uses cascaded multi-resolution radar inputs [16], and transformer-based models like T-FFTRadNet operating on raw ADC signals [8]. Researchers have also addressed efficient online detection using recurrent CNNs [4] and sparse networks like SparseRadNet [36]. More recently, methods including RTNH+ [11] and SpikingRTNH [23] have further advanced 4D radar object detection through improved pre-processing strategies and spiking neural network architectures. Several datasets support this line of research. RADIal [25] provides ADC signals, range-azimuth maps, and point clouds with segmentation annotations. Other datasets, including VoD [24], TJ4DRadSet [53], Dual Radar [48], MAN TruckScenes [5], and OmniHD-Scenes [54], primarily offer radar point clouds. In contrast, our work builds upon the K-Radar dataset [21], which is currently the only public 4D radar tensor dataset featuring multiple object types across diverse environments alongside corresponding 3D bounding box annotations. This unique combination enables a comprehensive evaluation of tensor-based 3D detection methods.
@@ III Method
We propose WRCFormer, a novel architecture designed for robust sensor fusion that processes inputs from a monocular camera and a 4D radar tensor as shown in Fig. 3. The framework begins with a parallel feature extraction pipeline, in which an Image Encoder and a Feature Pyramid Network (FPN) generate multi-scale visual features. Meanwhile, to reduce the resolution mismatch between image and radar data, we adopt a 4D radar tensor (cube) and decompose it into Elevation-Azimuth (EA) and Range-Azimuth (RA) representations. These radar maps are fed into specialized encoders and an FPN based on our proposed Wavelet Attention Mixture-of-Experts (WA-MoE) module, which decomposes them into complementary high- and low-frequency components and adaptively fuses them. To efficiently integrate multi-view features from both modalities while addressing the semantic–geometric misalignment often overlooked in existing fusion strategies, we introduce a Geometry-guided Progressive Fusion (GPF) module as the core fusion component of WRCFormer. GPF performs a two-stage coarse-to-fine fusion: In the first stage, Geometry-driven Semantic Alignment (GSA) aligns semantic image features with geometric EA radar maps via fast cross-attention. In the second stage, Range-aware Geometric Refinement (RGR) further refines the aligned visual features using deformable attention queries guided by the RA radar map. Finally, the fused features are forwarded to a detection head that applies iterative refinement to progressively improve predictions and produce the final 3D detection results.
@@ III-C Wavelet Attention Mixture-of-Experts
To better capture this structure, we design a wavelet-based neural network for frequency-domain decomposition. Among various wavelet bases, we adopt the Haar wavelet due to its computational simplicity, orthogonality, and suitability for representing piecewise-constant signals commonly encountered in radar amplitude maps. The Haar transform also aligns with the binary and multi-resolution nature of digital signal processing, making it efficient for real-time deployment. Our design is further motivated by the following aspects that collectively address the limitations of conventional convolution-based FPNs in radar tensor processing: (1) Noise components predominantly concentrate in the high-frequency sub-bands, whereas reflection peaks remain preserved in the low-frequency subbands. This property naturally facilitates target-clutter separation; (2) Radar sensing fundamentally operates in the frequency domain via Doppler and delay sampling, and wavelet tranform aligns well with the underlying radar scattering mechanism; (3) The introduction of a gated MoE is functionally equivalent to an adaptive filter bank, where the optimal spectral weights are estimated online for each channel and spatial location. This design enables data-driven spectrum shaping and offers greater flexibility than manually designed band-pass filters, while providing finer granularity than statically weighted Wavelet-CNNs; (4) For visual inputs, the high-frequency components of wavelets explicitly capture object edges, which further facilitate cross-modal alignment during the subsequent fusion process.
@@
where $\oplus$ denotes concatenation operation and $F_{1}$ is the first-order wavelet domain encoding result. This branch directly extracts spatial correlations within each frequency sub-band from the first-level wavelet decomposition. Consequently, it captures explicit and prominent spatio-frequency patterns, serving as a shallow feature extractor in the wavelet domain.
@@
where $\mathtt{g_{i}}(F_{\text{fused}})$ represents the sparse gating function, which dynamically computes a weighted combination of experts $\mathtt{E_{i}(\cdot)}$ for an input. Each expert is implemented as a lightweight Feed-Forward Network (FFN) using the depthwise-separable convolutions and specializes in different patterns.
@@ III-D Geometry-guided Progressive Fusion
To address the computational and alignment challenges in radar-camera fusion, we propose a Geometry-guided Progressive Fusion (GPF) module. GPF is a two-stage, coarse-to-fine fusion strategy designed to efficiently integrate multi-view radar and image features through structured geometric guidance, which includes Geometry-driven Semantic Alignment (GSA, Stage 1) and Range-aware Geometric Refinement (RGR, Stage 2).
@@
After that, we apply adaptive max-pooling to the EA query $Q\in\mathbb{R}^{N\times d}$ to obtain a compact representation $A\in\mathbb{R}^{n\times d}$, where $n\ll N$. Compared with adaptive average pooling, adaptive max pooling preserves the strongest signal responses in the EA features while effectively suppressing background interference, which preserves the most discriminative geometric responses while significantly reducing the sequence length.
@@
Table II: Overall comparison of average precision (AP) for 3D object detection on K-Radar (v1.0 version) with a threshold of 0.3.
@@ IV-B Quantitative Results
Overall Performances (v1.0 version). WRCFormer achieves the highest Total mAP3D of 58.7 (Table 1), outperforming camera-LiDAR fusion methods. This gain stems from our WA-MoE module and GPF mechanism, which enhance sparse radar representations and enable robust cross-modal alignment. In clear conditions (Normal/Overcast), camera-LiDAR fusion excels, benefiting from dense geometry. Conversely, in adverse weather (Fog, Rain, Sleet, Snow), our camera-radar fusion significantly outperforms camera-LiDAR alternatives, validating the effectiveness of wavelet-based frequency decomposition and geometry-guided fusion under sensor degradation. WRCFormer also surpasses other radar-camera methods (DPFT, EchoFusion) across most conditions, with a minor shortfall only against EchoFusion in Overcast weather, likely due to architectural differences in clear-sky feature extraction. The consistent advantage in challenging scenarios demonstrates the strength of our joint spatial-frequency encoding and progressive geometric alignment.
@@ V Conclusion
We introduced WRCFormer, a robust 3D object detector that effectively fused visual features with raw 4D radar tensors. The proposed method introduced two key innovations: a Wavelet Attention Mixture-of-Experts module that captures joint spatial-frequency representations to enhance sparse radar signatures, and a two-stage Geometry-guided Progressive Fusion mechanism that enables efficient and semantically-aligned cross-modal fusion integration. Extensive experiments on the K-Radar dataset proved that WRCFormer set a new state-of-the-art performance, with particularly strong robustness under adverse weather conditions where camera-based perception often degrades. These results underscored the significant advantage of directly leveraging rich 4D radar tensor, paired with specialized neural architectures, for building reliable, all-weather perception systems in autonomous driving.